OpenStudy (anonymous):

how would i do 10P4 by hand?

OpenStudy (solomonzelman):

\(\large\color{midnightblue}{ \rm nPr~=~n!~/~(n-r)! }\)

OpenStudy (anonymous):

thanks man. what about nCr?

OpenStudy (solomonzelman):

\(\large\color{midnightblue}{ \rm 10~P~4~=~10!~/~(10-4)!= }\) \(\large\color{midnightblue}{ \rm 10!~/~6! }\) \(\large\color{midnightblue}{ \rm (6! \times 7 \times 8 \times 9 \times 10)~/~6! }\)

OpenStudy (solomonzelman):

The rest you can do yourself. What about nCr, \(\large\color{midnightblue}{ \rm nCr=n!~~/~~r!(n-r)! }\)

OpenStudy (anonymous):

awesome. thanks so much.

OpenStudy (solomonzelman):

Anytime :)

OpenStudy (solomonzelman):

` Also, for a future reference: ` ` ` `you can find any section by: ` `1) clicking "more subjects" (the top left corner, towards the center) ` `2) typing in the name of the section (name of a section can be ` `"history" Mathematics and other section for each unique area of ` `studying). ` ` ` `You can FAN someone by: ` `1) clicking on the user's name ` `2) clicking "fan" on the bottom right corner of user's box. ` `( you can also write a testimonial by clicking "add testimonial" and ` `typing something in testimonial box) ` ` ` ` ` ` ` `you can mention someone in your question (or in any other question) ` `by writing ` @ `in front of their name, just like I would mention you... ` `LOOK ! .... ` @fenton `(they would get a notification that a user ` `mentioned them in a reply) ` ` ` ` ` `you can message someone by: ` `1) clicking on the someone's username ` `2) clicking "Send Message" ` `3) type your message press submit, and wait till the messages ` `submits. ` ` ` ` ` `If you have any questions you can message me or anyone else on this ` ` website. We will be happy to respond ! `